OpenStudy (anonymous):

factor: see below:

OpenStudy (anonymous):

\[81z ^{2}-4k ^{2}\]

jimthompson5910 (jim_thompson5910):

hint: a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

please show steps because I am confused. Thank you

OpenStudy (anonymous):

9^2?

OpenStudy (anonymous):

9^2-2^2

jimthompson5910 (jim_thompson5910):

81z ^2 = (9z)^2 and 4k^2 = (2k)^2

jimthompson5910 (jim_thompson5910):

so you really have (9z)^2 - (2k)^2 now use the rule given above

OpenStudy (tpaulus):

It is a difference of squares a^2 = 81 b^2 = 4 (aZ+bK)(aZ-bK)

OpenStudy (anonymous):

ok thanks a lot

OpenStudy (anonymous):

how do you know if positive or negative in answer?

OpenStudy (anonymous):

and which is positive or negative

OpenStudy (anonymous):

if the problem is negative. means every factored answer has a one negative and one positive?

OpenStudy (anonymous):

answer: \[(9z ^{2}-2k ^{2})(9z ^{2}+2k ^{2})\]

jimthompson5910 (jim_thompson5910):

not sure what you mean, but you'll have a positive portion and a negative portion

jimthompson5910 (jim_thompson5910):

no it should be (9z - 2k)(9z + 2k)

jimthompson5910 (jim_thompson5910):

using the difference of squares rule

OpenStudy (anonymous):

ok I think I got it now

jimthompson5910 (jim_thompson5910):

ok great